Find two consecutive odd numbers such that the sum of three-sevenths of the first number and one-third of the second number is equal to thirty-eight.
Answer by checkley71(8403)

LET X & X+2 BE THE 2 CONSECUTIVE ODD NUMBERS.
3X/7+(X+2)/3=38
[3*3X+7(X+2)]/3*7=38
[9X+7X+14]/21=38 CROSS MULTIPLY
16X+14=21*38
16X=798-14
16X=784
X=784/16
X=49 THE SMALLER ODD NUMBER
49+2=51 THE LARGER ODD NUMBER.
PROOF
3*49/7+51/3=38
3*7+17=38
21+17=38
38=38